If an object has a mass of 47 kg and it is moved 27 meters in 60 seconds, how

much power was used



1
Expert's answer
2021-07-22T10:02:58-0400

By the definition of the power, we have:


"P=Fv=mg\dfrac{d}{t},""P=47\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot\dfrac{27\ m}{60\ s}=207\ W."